If the larger of two integers, whose sum is 88, is divided by the smaller, the quotient is 5
and the remainder is 10. What are the two numbers ?

Answer by Boreal(15235)

larger integer is x
smaller is 88-x
x/88-x =5 remainder 10/88-x
(x/88-x)=5+(10/88-x)
common denominator of 88-x
(x/88-x)=(440-5x+10)/88-x
Therefore, x= -5x+450, since the denominators are equal.
6x=450
x=75
The larger integer is 75
The smaller is 13
They add to 88,
Their quotient is 5 remainder 10
The remainder is another way of showing a fraction, whose numerator is the remainder, the denominator the divisor.